# Faster Neighborhood Attention: Reducing the $\mathcal{O}(n^{2})$ Cost of Self Attention at the Threadblock Level

### v2

@@ 4 Experiments
We further present a breakdown of our benchmarks in Tab. 3 and Tab. 4, where we group cases by window size, which is the primary contributor to difference in FLOPs when considering similar input sizes. We report the average, minimum, and maximum improvement observed from switching from naive to GEMM-based, naive to fused, and GEMM-based to fused kernels. GEMM-based kernels exhibit strong performance compared to both naive and fused kernels in full precision, where fused kernels only have a very minimal edge on average. GEMM-based kernels also outperform naive kernels in half precision, especially in cases where tiled kernels are not available. While tiled kernels are sometimes the better choice, we note that they simply cannot generalize to all problem sizes as our GEMM-based kernels can, nor are they easily extensible. In addition, our auto-tuner can be extended to include BMM-style implementations as well as fused kernels, and pick the best between GEMM-based and naive (naive or tiled) according to problem size, hardware, and other such performance determining factors.
Window GEMM over naive Fused over naive Fused over GEMM
size Average Min Max Average Min Max Average Min Max
1-dimensional neighborhood attention
9 $\boldsymbol{\uparrow}$ 199 % $\boldsymbol{\downarrow}$59 % $\boldsymbol{\uparrow}$ 673 % $\boldsymbol{\uparrow}$ 705 % $\boldsymbol{\uparrow}$ 17 % $\boldsymbol{\uparrow}$ 1660 % $\boldsymbol{\uparrow}$ 178 % $\boldsymbol{\uparrow}$ 44 % $\boldsymbol{\uparrow}$ 300 %
15 $\boldsymbol{\uparrow}$ 258 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 929 % $\boldsymbol{\uparrow}$ 844 % $\boldsymbol{\uparrow}$ 57 % $\boldsymbol{\uparrow}$ 2000 % $\boldsymbol{\uparrow}$ 181 % $\boldsymbol{\uparrow}$ 49 % $\boldsymbol{\uparrow}$ 280 %
31 $\boldsymbol{\uparrow}$ 412 % $\boldsymbol{\downarrow}$12 % $\boldsymbol{\uparrow}$ 1567 % $\boldsymbol{\uparrow}$ 1217 % $\boldsymbol{\uparrow}$ 150 % $\boldsymbol{\uparrow}$ 3206 % $\boldsymbol{\uparrow}$ 183 % $\boldsymbol{\uparrow}$ 62 % $\boldsymbol{\uparrow}$ 280 %
45 $\boldsymbol{\uparrow}$ 356 % $\boldsymbol{\downarrow}$22 % $\boldsymbol{\uparrow}$ 1429 % $\boldsymbol{\uparrow}$ 1100 % $\boldsymbol{\uparrow}$ 133 % $\boldsymbol{\uparrow}$ 3150 % $\boldsymbol{\uparrow}$ 190 % $\boldsymbol{\uparrow}$ 77 % $\boldsymbol{\uparrow}$ 320 %
63 $\boldsymbol{\uparrow}$ 427 % $\boldsymbol{\downarrow}$28 % $\boldsymbol{\uparrow}$ 1697 % $\boldsymbol{\uparrow}$ 1275 % $\boldsymbol{\uparrow}$ 117 % $\boldsymbol{\uparrow}$ 3981 % $\boldsymbol{\uparrow}$ 191 % $\boldsymbol{\uparrow}$ 50 % $\boldsymbol{\uparrow}$ 320 %
127 $\boldsymbol{\uparrow}$ 545 % $\boldsymbol{\downarrow}$21 % $\boldsymbol{\uparrow}$ 2430 % $\boldsymbol{\uparrow}$ 1497 % $\boldsymbol{\uparrow}$ 73 % $\boldsymbol{\uparrow}$ 7494 % $\boldsymbol{\uparrow}$ 167 % $\boldsymbol{\uparrow}$ 62 % $\boldsymbol{\uparrow}$ 360 %
255 $\boldsymbol{\uparrow}$ 779 % $\boldsymbol{\downarrow}$11 % $\boldsymbol{\uparrow}$ 2598 % $\boldsymbol{\uparrow}$ 2231 % $\boldsymbol{\uparrow}$ 164 % $\boldsymbol{\uparrow}$ 8797 % $\boldsymbol{\uparrow}$ 171 % $\boldsymbol{\uparrow}$ 94 % $\boldsymbol{\uparrow}$ 291 %
511 $\boldsymbol{\uparrow}$ 1029 % $\boldsymbol{\uparrow}$ 23 % $\boldsymbol{\uparrow}$ 2781 % $\boldsymbol{\uparrow}$ 3369 % $\boldsymbol{\uparrow}$ 247 % $\boldsymbol{\uparrow}$ 10649 % $\boldsymbol{\uparrow}$ 207 % $\boldsymbol{\uparrow}$ 104 % $\boldsymbol{\uparrow}$ 344 %
767 $\boldsymbol{\uparrow}$ 1193 % $\boldsymbol{\uparrow}$ 59 % $\boldsymbol{\uparrow}$ 2473 % $\boldsymbol{\uparrow}$ 4278 % $\boldsymbol{\uparrow}$ 243 % $\boldsymbol{\uparrow}$ 10226 % $\boldsymbol{\uparrow}$ 241 % $\boldsymbol{\uparrow}$ 111 % $\boldsymbol{\uparrow}$ 386 %
1023 $\boldsymbol{\uparrow}$ 1377 % $\boldsymbol{\uparrow}$ 101 % $\boldsymbol{\uparrow}$ 2783 % $\boldsymbol{\uparrow}$ 5030 % $\boldsymbol{\uparrow}$ 444 % $\boldsymbol{\uparrow}$ 11082 % $\boldsymbol{\uparrow}$ 255 % $\boldsymbol{\uparrow}$ 155 % $\boldsymbol{\uparrow}$ 359 %
2047 $\boldsymbol{\uparrow}$ 1342 % $\boldsymbol{\uparrow}$ 181 % $\boldsymbol{\uparrow}$ 2665 % $\boldsymbol{\uparrow}$ 5721 % $\boldsymbol{\uparrow}$ 765 % $\boldsymbol{\uparrow}$ 11540 % $\boldsymbol{\uparrow}$ 322 % $\boldsymbol{\uparrow}$ 159 % $\boldsymbol{\uparrow}$ 444 %
4095 $\boldsymbol{\uparrow}$ 1434 % $\boldsymbol{\uparrow}$ 203 % $\boldsymbol{\uparrow}$ 2787 % $\boldsymbol{\uparrow}$ 6027 % $\boldsymbol{\uparrow}$ 1223 % $\boldsymbol{\uparrow}$ 11093 % $\boldsymbol{\uparrow}$ 323 % $\boldsymbol{\uparrow}$ 210 % $\boldsymbol{\uparrow}$ 453 %
All $\boldsymbol{\uparrow}$ 496 % $\boldsymbol{\downarrow}$59 % $\boldsymbol{\uparrow}$ 2787 % $\boldsymbol{\uparrow}$ 1607 % $\boldsymbol{\uparrow}$ 17 % $\boldsymbol{\uparrow}$ 11540 % $\boldsymbol{\uparrow}$ 190 % $\boldsymbol{\uparrow}$ 44 % $\boldsymbol{\uparrow}$ 453 %
2-dimensional neighborhood attention
3 × 3 $\boldsymbol{\uparrow}$ 27 % $\boldsymbol{\downarrow}$68 % $\boldsymbol{\uparrow}$ 258 % $\boldsymbol{\uparrow}$ 203 % $\boldsymbol{\downarrow}$43 % $\boldsymbol{\uparrow}$ 900 % $\boldsymbol{\uparrow}$ 164 % $\boldsymbol{\uparrow}$ 18 % $\boldsymbol{\uparrow}$ 764 %
5 × 5 $\boldsymbol{\uparrow}$ 101 % $\boldsymbol{\downarrow}$68 % $\boldsymbol{\uparrow}$ 554 % $\boldsymbol{\uparrow}$ 420 % $\boldsymbol{\downarrow}$25 % $\boldsymbol{\uparrow}$ 1990 % $\boldsymbol{\uparrow}$ 180 % $\boldsymbol{\uparrow}$ 24 % $\boldsymbol{\uparrow}$ 862 %
7 × 7 $\boldsymbol{\uparrow}$ 108 % $\boldsymbol{\downarrow}$61 % $\boldsymbol{\uparrow}$ 570 % $\boldsymbol{\uparrow}$ 509 % $\boldsymbol{\downarrow}$40 % $\boldsymbol{\uparrow}$ 3086 % $\boldsymbol{\uparrow}$ 211 % $\boldsymbol{\uparrow}$ 15 % $\boldsymbol{\uparrow}$ 1032 %
9 × 9 $\boldsymbol{\uparrow}$ 129 % $\boldsymbol{\downarrow}$52 % $\boldsymbol{\uparrow}$ 444 % $\boldsymbol{\uparrow}$ 505 % $\boldsymbol{\downarrow}$20 % $\boldsymbol{\uparrow}$ 1772 % $\boldsymbol{\uparrow}$ 163 % $\boldsymbol{\uparrow}$ 38 % $\boldsymbol{\uparrow}$ 434 %
11 × 11 $\boldsymbol{\uparrow}$ 132 % $\boldsymbol{\downarrow}$52 % $\boldsymbol{\uparrow}$ 531 % $\boldsymbol{\uparrow}$ 520 % $\boldsymbol{\downarrow}$27 % $\boldsymbol{\uparrow}$ 2298 % $\boldsymbol{\uparrow}$ 167 % $\boldsymbol{\uparrow}$ 28 % $\boldsymbol{\uparrow}$ 437 %
13 × 13 $\boldsymbol{\uparrow}$ 121 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 504 % $\boldsymbol{\uparrow}$ 697 % $\boldsymbol{\downarrow}$30 % $\boldsymbol{\uparrow}$ 3037 % $\boldsymbol{\uparrow}$ 252 % $\boldsymbol{\uparrow}$ 37 % $\boldsymbol{\uparrow}$ 593 %
15 × 15 $\boldsymbol{\uparrow}$ 207 % $\boldsymbol{\downarrow}$8 % $\boldsymbol{\uparrow}$ 740 % $\boldsymbol{\uparrow}$ 1082 % $\boldsymbol{\uparrow}$ 79 % $\boldsymbol{\uparrow}$ 5086 % $\boldsymbol{\uparrow}$ 269 % $\boldsymbol{\uparrow}$ 50 % $\boldsymbol{\uparrow}$ 618 %
17 × 17 $\boldsymbol{\uparrow}$ 251 % $\boldsymbol{\uparrow}$ 20 % $\boldsymbol{\uparrow}$ 732 % $\boldsymbol{\uparrow}$ 883 % $\boldsymbol{\uparrow}$ 106 % $\boldsymbol{\uparrow}$ 2384 % $\boldsymbol{\uparrow}$ 178 % $\boldsymbol{\uparrow}$ 53 % $\boldsymbol{\uparrow}$ 284 %
33 × 33 $\boldsymbol{\uparrow}$ 178 % $\boldsymbol{\downarrow}$13 % $\boldsymbol{\uparrow}$ 372 % $\boldsymbol{\uparrow}$ 1686 % $\boldsymbol{\uparrow}$ 469 % $\boldsymbol{\uparrow}$ 3006 % $\boldsymbol{\uparrow}$ 558 % $\boldsymbol{\uparrow}$ 429 % $\boldsymbol{\uparrow}$ 675 %
41 × 41 $\boldsymbol{\uparrow}$ 208 % $\boldsymbol{\downarrow}$8 % $\boldsymbol{\uparrow}$ 426 % $\boldsymbol{\uparrow}$ 1969 % $\boldsymbol{\uparrow}$ 588 % $\boldsymbol{\uparrow}$ 3421 % $\boldsymbol{\uparrow}$ 591 % $\boldsymbol{\uparrow}$ 400 % $\boldsymbol{\uparrow}$ 724 %
49 × 49 $\boldsymbol{\uparrow}$ 210 % $\boldsymbol{\downarrow}$12 % $\boldsymbol{\uparrow}$ 427 % $\boldsymbol{\uparrow}$ 2357 % $\boldsymbol{\uparrow}$ 753 % $\boldsymbol{\uparrow}$ 4081 % $\boldsymbol{\uparrow}$ 718 % $\boldsymbol{\uparrow}$ 468 % $\boldsymbol{\uparrow}$ 874 %
63 × 63 $\boldsymbol{\uparrow}$ 246 % $\boldsymbol{\uparrow}$ 5 % $\boldsymbol{\uparrow}$ 490 % $\boldsymbol{\uparrow}$ 3165 % $\boldsymbol{\uparrow}$ 922 % $\boldsymbol{\uparrow}$ 5510 % $\boldsymbol{\uparrow}$ 872 % $\boldsymbol{\uparrow}$ 419 % $\boldsymbol{\uparrow}$ 1096 %
All $\boldsymbol{\uparrow}$ 113 % $\boldsymbol{\downarrow}$68 % $\boldsymbol{\uparrow}$ 740 % $\boldsymbol{\uparrow}$ 581 % $\boldsymbol{\downarrow}$43 % $\boldsymbol{\uparrow}$ 5510 % $\boldsymbol{\uparrow}$ 213 % $\boldsymbol{\uparrow}$ 15 % $\boldsymbol{\uparrow}$ 1096 %
3-dimensional neighborhood attention
3 × 3 × 3 - - - $\boldsymbol{\uparrow}$ 341 % 0 % $\boldsymbol{\uparrow}$ 1651 % - - -
5 × 5 × 5 - - - $\boldsymbol{\uparrow}$ 318 % $\boldsymbol{\downarrow}$27 % $\boldsymbol{\uparrow}$ 1932 % - - -
7 × 7 × 7 - - - $\boldsymbol{\uparrow}$ 667 % $\boldsymbol{\uparrow}$ 42 % $\boldsymbol{\uparrow}$ 3583 % - - -
9 × 9 × 9 - - - $\boldsymbol{\uparrow}$ 602 % $\boldsymbol{\uparrow}$ 98 % $\boldsymbol{\uparrow}$ 1550 % - - -
11 × 11 × 11 - - - $\boldsymbol{\uparrow}$ 878 % $\boldsymbol{\uparrow}$ 160 % $\boldsymbol{\uparrow}$ 2229 % - - -
13 × 13 × 13 - - - $\boldsymbol{\uparrow}$ 1208 % $\boldsymbol{\uparrow}$ 237 % $\boldsymbol{\uparrow}$ 3130 % - - -
15 × 15 × 15 - - - $\boldsymbol{\uparrow}$ 1600 % $\boldsymbol{\uparrow}$ 241 % $\boldsymbol{\uparrow}$ 4376 % - - -
17 × 17 × 17 - - - $\boldsymbol{\uparrow}$ 822 % $\boldsymbol{\uparrow}$ 303 % $\boldsymbol{\uparrow}$ 1473 % - - -
All - - - $\boldsymbol{\uparrow}$ 552 % $\boldsymbol{\downarrow}$27 % $\boldsymbol{\uparrow}$ 4376 % - - -
Table 3: Breakdown of improvement over naive kernels (FP16). Both GEMM-based and fused NA improve the baseline naive kernels on average. However, as illustrated by the minimum improvement, there are many cases in which naive kernels may be preferable to GEMM-based, which fails to effectively hide data prefetching latency. Fused NA on the other hand manages to improve upon most cases, with the exception of many cases that target the tiled kernels in NATTEN (2-dimensional, with window size up to 13 × 13.) Very few cases ($\approx$ 0.3 %) in 3-dimensional problems where fused NA can’t always outperform naive.
Window GEMM over naive Fused over naive Fused over GEMM
size Average Min Max Average Min Max Average Min Max
1-dimensional neighborhood attention
9 $\boldsymbol{\uparrow}$ 473 % $\boldsymbol{\downarrow}$40 % $\boldsymbol{\uparrow}$ 1046 % $\boldsymbol{\uparrow}$ 639 % $\boldsymbol{\downarrow}$10 % $\boldsymbol{\uparrow}$ 1682 % $\boldsymbol{\uparrow}$ 29 % $\boldsymbol{\downarrow}$48 % $\boldsymbol{\uparrow}$ 109 %
15 $\boldsymbol{\uparrow}$ 584 % $\boldsymbol{\downarrow}$33 % $\boldsymbol{\uparrow}$ 1462 % $\boldsymbol{\uparrow}$ 739 % $\boldsymbol{\uparrow}$ 11 % $\boldsymbol{\uparrow}$ 1718 % $\boldsymbol{\uparrow}$ 27 % $\boldsymbol{\downarrow}$45 % $\boldsymbol{\uparrow}$ 100 %
31 $\boldsymbol{\uparrow}$ 829 % $\boldsymbol{\uparrow}$ 106 % $\boldsymbol{\uparrow}$ 2192 % $\boldsymbol{\uparrow}$ 1002 % $\boldsymbol{\uparrow}$ 230 % $\boldsymbol{\uparrow}$ 2247 % $\boldsymbol{\uparrow}$ 26 % $\boldsymbol{\downarrow}$38 % $\boldsymbol{\uparrow}$ 118 %
45 $\boldsymbol{\uparrow}$ 678 % $\boldsymbol{\uparrow}$ 100 % $\boldsymbol{\uparrow}$ 1925 % $\boldsymbol{\uparrow}$ 803 % $\boldsymbol{\uparrow}$ 182 % $\boldsymbol{\uparrow}$ 1981 % $\boldsymbol{\uparrow}$ 22 % $\boldsymbol{\downarrow}$37 % $\boldsymbol{\uparrow}$ 108 %
63 $\boldsymbol{\uparrow}$ 749 % $\boldsymbol{\uparrow}$ 50 % $\boldsymbol{\uparrow}$ 2118 % $\boldsymbol{\uparrow}$ 906 % $\boldsymbol{\uparrow}$ 109 % $\boldsymbol{\uparrow}$ 2567 % $\boldsymbol{\uparrow}$ 24 % $\boldsymbol{\downarrow}$35 % $\boldsymbol{\uparrow}$ 117 %
127 $\boldsymbol{\uparrow}$ 874 % $\boldsymbol{\uparrow}$ 26 % $\boldsymbol{\uparrow}$ 3139 % $\boldsymbol{\uparrow}$ 934 % $\boldsymbol{\uparrow}$ 26 % $\boldsymbol{\uparrow}$ 4179 % $\boldsymbol{\uparrow}$ 6 % $\boldsymbol{\downarrow}$39 % $\boldsymbol{\uparrow}$ 95 %
255 $\boldsymbol{\uparrow}$ 1190 % $\boldsymbol{\uparrow}$ 52 % $\boldsymbol{\uparrow}$ 3237 % $\boldsymbol{\uparrow}$ 1286 % $\boldsymbol{\uparrow}$ 45 % $\boldsymbol{\uparrow}$ 4767 % 0 % $\boldsymbol{\downarrow}$38 % $\boldsymbol{\uparrow}$ 69 %
511 $\boldsymbol{\uparrow}$ 1681 % $\boldsymbol{\uparrow}$ 196 % $\boldsymbol{\uparrow}$ 3404 % $\boldsymbol{\uparrow}$ 1894 % $\boldsymbol{\uparrow}$ 86 % $\boldsymbol{\uparrow}$ 4794 % $\boldsymbol{\uparrow}$ 5 % $\boldsymbol{\downarrow}$50 % $\boldsymbol{\uparrow}$ 67 %
767 $\boldsymbol{\uparrow}$ 2016 % $\boldsymbol{\uparrow}$ 475 % $\boldsymbol{\uparrow}$ 3276 % $\boldsymbol{\uparrow}$ 2429 % $\boldsymbol{\uparrow}$ 229 % $\boldsymbol{\uparrow}$ 4623 % $\boldsymbol{\uparrow}$ 16 % $\boldsymbol{\downarrow}$51 % $\boldsymbol{\uparrow}$ 75 %
1023 $\boldsymbol{\uparrow}$ 2209 % $\boldsymbol{\uparrow}$ 649 % $\boldsymbol{\uparrow}$ 3488 % $\boldsymbol{\uparrow}$ 2638 % $\boldsymbol{\uparrow}$ 288 % $\boldsymbol{\uparrow}$ 4668 % $\boldsymbol{\uparrow}$ 16 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 65 %
2047 $\boldsymbol{\uparrow}$ 2265 % $\boldsymbol{\uparrow}$ 888 % $\boldsymbol{\uparrow}$ 3458 % $\boldsymbol{\uparrow}$ 2923 % $\boldsymbol{\uparrow}$ 618 % $\boldsymbol{\uparrow}$ 4656 % $\boldsymbol{\uparrow}$ 28 % $\boldsymbol{\downarrow}$34 % $\boldsymbol{\uparrow}$ 74 %
4095 $\boldsymbol{\uparrow}$ 2319 % $\boldsymbol{\uparrow}$ 1125 % $\boldsymbol{\uparrow}$ 3546 % $\boldsymbol{\uparrow}$ 3009 % $\boldsymbol{\uparrow}$ 1056 % $\boldsymbol{\uparrow}$ 4679 % $\boldsymbol{\uparrow}$ 29 % $\boldsymbol{\downarrow}$9 % $\boldsymbol{\uparrow}$ 75 %
All $\boldsymbol{\uparrow}$ 895 % $\boldsymbol{\downarrow}$40 % $\boldsymbol{\uparrow}$ 3546 % $\boldsymbol{\uparrow}$ 1072 % $\boldsymbol{\downarrow}$10 % $\boldsymbol{\uparrow}$ 4794 % $\boldsymbol{\uparrow}$ 21 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 118 %
2-dimensional neighborhood attention
3 × 3 $\boldsymbol{\uparrow}$ 135 % $\boldsymbol{\downarrow}$58 % $\boldsymbol{\uparrow}$ 607 % $\boldsymbol{\uparrow}$ 161 % $\boldsymbol{\downarrow}$65 % $\boldsymbol{\uparrow}$ 810 % $\boldsymbol{\uparrow}$ 29 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 301 %
5 × 5 $\boldsymbol{\uparrow}$ 260 % $\boldsymbol{\downarrow}$52 % $\boldsymbol{\uparrow}$ 1021 % $\boldsymbol{\uparrow}$ 290 % $\boldsymbol{\downarrow}$62 % $\boldsymbol{\uparrow}$ 1301 % $\boldsymbol{\uparrow}$ 28 % $\boldsymbol{\downarrow}$51 % $\boldsymbol{\uparrow}$ 348 %
7 × 7 $\boldsymbol{\uparrow}$ 246 % $\boldsymbol{\downarrow}$47 % $\boldsymbol{\uparrow}$ 1006 % $\boldsymbol{\uparrow}$ 331 % $\boldsymbol{\downarrow}$67 % $\boldsymbol{\uparrow}$ 2005 % $\boldsymbol{\uparrow}$ 45 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 436 %
9 × 9 $\boldsymbol{\uparrow}$ 238 % $\boldsymbol{\downarrow}$32 % $\boldsymbol{\uparrow}$ 648 % $\boldsymbol{\uparrow}$ 303 % $\boldsymbol{\downarrow}$54 % $\boldsymbol{\uparrow}$ 1052 % $\boldsymbol{\uparrow}$ 26 % $\boldsymbol{\downarrow}$46 % $\boldsymbol{\uparrow}$ 156 %
11 × 11 $\boldsymbol{\uparrow}$ 244 % $\boldsymbol{\downarrow}$29 % $\boldsymbol{\uparrow}$ 856 % $\boldsymbol{\uparrow}$ 294 % $\boldsymbol{\downarrow}$61 % $\boldsymbol{\uparrow}$ 1459 % $\boldsymbol{\uparrow}$ 20 % $\boldsymbol{\downarrow}$54 % $\boldsymbol{\uparrow}$ 144 %
13 × 13 $\boldsymbol{\uparrow}$ 220 % $\boldsymbol{\downarrow}$27 % $\boldsymbol{\uparrow}$ 789 % $\boldsymbol{\uparrow}$ 393 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 1835 % $\boldsymbol{\uparrow}$ 59 % $\boldsymbol{\downarrow}$50 % $\boldsymbol{\uparrow}$ 217 %
15 × 15 $\boldsymbol{\uparrow}$ 401 % $\boldsymbol{\downarrow}$10 % $\boldsymbol{\uparrow}$ 1175 % $\boldsymbol{\uparrow}$ 698 % 0 % $\boldsymbol{\uparrow}$ 2845 % $\boldsymbol{\uparrow}$ 56 % $\boldsymbol{\downarrow}$45 % $\boldsymbol{\uparrow}$ 240 %
17 × 17 $\boldsymbol{\uparrow}$ 510 % $\boldsymbol{\uparrow}$ 90 % $\boldsymbol{\uparrow}$ 1246 % $\boldsymbol{\uparrow}$ 565 % $\boldsymbol{\uparrow}$ 23 % $\boldsymbol{\uparrow}$ 1347 % $\boldsymbol{\uparrow}$ 10 % $\boldsymbol{\downarrow}$45 % $\boldsymbol{\uparrow}$ 73 %
33 × 33 $\boldsymbol{\uparrow}$ 860 % $\boldsymbol{\uparrow}$ 308 % $\boldsymbol{\uparrow}$ 1305 % $\boldsymbol{\uparrow}$ 1209 % $\boldsymbol{\uparrow}$ 507 % $\boldsymbol{\uparrow}$ 1746 % $\boldsymbol{\uparrow}$ 40 % $\boldsymbol{\downarrow}$5 % $\boldsymbol{\uparrow}$ 81 %
41 × 41 $\boldsymbol{\uparrow}$ 966 % $\boldsymbol{\uparrow}$ 356 % $\boldsymbol{\uparrow}$ 1463 % $\boldsymbol{\uparrow}$ 1418 % $\boldsymbol{\uparrow}$ 541 % $\boldsymbol{\uparrow}$ 2068 % $\boldsymbol{\uparrow}$ 46 % $\boldsymbol{\downarrow}$26 % $\boldsymbol{\uparrow}$ 90 %
49 × 49 $\boldsymbol{\uparrow}$ 906 % $\boldsymbol{\uparrow}$ 345 % $\boldsymbol{\uparrow}$ 1299 % $\boldsymbol{\uparrow}$ 1719 % $\boldsymbol{\uparrow}$ 673 % $\boldsymbol{\uparrow}$ 2524 % $\boldsymbol{\uparrow}$ 86 % $\boldsymbol{\downarrow}$9 % $\boldsymbol{\uparrow}$ 144 %
[…]

### v3

@@ 4 Experiments
We further present a breakdown of our benchmarks in Tab. 3, where we report the average, minimum, and maximum improvement observed from switching from naive to GEMM-based, naive to fused, and GEMM-based to fused kernels. GEMM-based kernels exhibit strong performance compared to both naive and fused kernels in full precision, where fused kernels only have a very minimal edge over unfused. GEMM-based kernels also outperform naive kernels in half precision, especially in cases where tiled kernels are not available. While the tiled kernels are sometimes the better choice, we note that they simply cannot generalize to all problem sizes as our GEMM-based kernels can, nor are they easily extensible.
Table 3: Forward pass benchmark breakdown. Both GEMM-based and fused NA improve the baseline naive kernels on average. However, there exist cases in which naive kernels may be preferable to GEMM-based in both FP16 and FP32, but naive is rarely a good choice in half precision where both naive and GEMM are more memory bandwidth bound than fused.
Dim GEMM over naive Fused over naive Fused over GEMM
Average Min Max Average Min Max Average Min Max
FP16
1-D $\boldsymbol{\uparrow}$ 548 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 3025 % $\boldsymbol{\uparrow}$ 1759 % $\boldsymbol{\uparrow}$ 60 % $\boldsymbol{\uparrow}$ 11885 % $\boldsymbol{\uparrow}$ 180 % $\boldsymbol{\uparrow}$ 71 % $\boldsymbol{\uparrow}$ 466 %
2-D $\boldsymbol{\uparrow}$ 193 % $\boldsymbol{\downarrow}$57 % $\boldsymbol{\uparrow}$ 862 % $\boldsymbol{\uparrow}$ 958 % 0 % $\boldsymbol{\uparrow}$ 7169 % $\boldsymbol{\uparrow}$ 257 % $\boldsymbol{\uparrow}$ 38 % $\boldsymbol{\uparrow}$ 1199 %
3-D - - - $\boldsymbol{\uparrow}$ 1135 % $\boldsymbol{\uparrow}$ 118 % $\boldsymbol{\uparrow}$ 5497 % - - -
FP32
1-D $\boldsymbol{\uparrow}$ 874 % $\boldsymbol{\downarrow}$31 % $\boldsymbol{\uparrow}$ 3565 % $\boldsymbol{\uparrow}$ 978 % $\boldsymbol{\uparrow}$ 13 % $\boldsymbol{\uparrow}$ 4419 % $\boldsymbol{\uparrow}$ 17 % $\boldsymbol{\downarrow}$54 % $\boldsymbol{\uparrow}$ 136 %
2-D $\boldsymbol{\uparrow}$ 386 % $\boldsymbol{\downarrow}$43 % $\boldsymbol{\uparrow}$ 1933 % $\boldsymbol{\uparrow}$ 564 % $\boldsymbol{\downarrow}$30 % $\boldsymbol{\uparrow}$ 4043 % $\boldsymbol{\uparrow}$ 43 % $\boldsymbol{\downarrow}$53 % $\boldsymbol{\uparrow}$ 451 %
3-D - - - $\boldsymbol{\uparrow}$ 712 % $\boldsymbol{\uparrow}$ 25 % $\boldsymbol{\uparrow}$ 3029 % - - -
5 Future work & Conclusion
In this work, we formulated the neighborhood attention problem, and by extension multi-dimensional sliding window attention, which are inherently GEMV problems, as GEMM/GETT problems. Through this finding, we implemented extensible GEMM-based and fused CUDA kernels that implement neighborhood attention, which can significantly improve upon existing kernels in the $\mathcal{N}ATTEN$ project. These kernels will not only speed up previously-proposed models based on neighborhood attention, but can also significantly enhance ongoing research efforts in this direction. In addition, our fused kernels are the most flexible in terms of parameterization, by supporting varying window sizes, dilation factors, and causal masking across different axes, which enable unique applications such as 3-D spatio-temporal attention with causal masking across time. They also enjoy a reduced memory footprint, and can avoid being bound by memory bandwidth at scale.
Future directions in this area include but are not limited to: support for Context Parallelism (CP), implementations using more efficient predication (i.e. with the Hopper TMA), extension to more modern architectures (warp-specialized kernels in Hopper and Blackwell), extension to other AI accelerators, and better auto-tuning (or alternatives involving graph compilation).
We’ve shown that multi-dimensional local attention can indeed serve as solutions for scaling future large-scale long-context architectures, when provided with suitable software infrastructure. We hope that this inspires more research into multi-dimensional attention, as deep learning systems continue to grow larger in both model and input size.
Acknowledgements. We would like to thank NVIDIA and members of the CUTLASS project, in particular Haicheng Wu, for his valuable feedback and comments which led to the creation of GEMM-based NA. We also thank Meta xFormers team for developing FMHA, which is what our fused neighborhood attention kernels are based on. A. H. thanks Michael Isaev, Aditya Kane, and Kai Wang for their feedback on the paper. A. H. also thanks Bing Xu, Hao Lu, Michael Iovine, and Terry Chen for the invaluable learning experience while interning at HippoML, which helped accelerate the timeline of this project. This research was supported in part by National Science Foundation under Award #2427478 - CAREER Program, and by National Science Foundation and the Institute of Education Sciences, U.S. Department of Education under Award #2229873 - National AI Institute for Exceptional Education. This project was also partially supported by cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at the Georgia Institute of Technology, Atlanta, Georgia, USA.
Appendix A Auto-tuner
GEMM kernels are, among other settings, parameterized by their tiling shape. Multi-dimensional variants (2-D and 3-D neighborhood attention) can also be parameterized by their fine-grained tile sizes, introduced by our formulation. As mentioned earlier, a GEMM with row tile size 64 can be reinterpreted as a number of 2-D and 3-D tiles (i.e. $x\times y$ for all positive integers $x$ and $y$ where $xy=64$, and $x\times y\times z$ for all positive integers $x$, $y$, and $z$ where $xyz=64$.) As a result, selecting tiling sizes based on factors such as problem size, hardware, and environment can further decrease achievable runtime. We therefore implement a very simple auto-tuning method as a proof of concept. Auto-tuning creates and maintains a cache for the lifetime of the application, which maps problems (defined by problem size, data type, and other such factors) to a tiling configuration. On a cache miss, the problem is benchmarked over a set of tiling configurations, and the best configuration gets cached.
While the auto-tuner can noticeably improve performance even further, we note that it is presently limited in the following:
  1. 1.
    Distributed training. auto-tuner context is limited to a single process, meaning jobs involving distributed training or inference will run the auto-tuner separately in each individual process. Aside from the possibility of different processes choosing different settings, which can slightly impact numerical determinism, this behavior is counter-intuitive. A more advanced auto-tuner would distribute possible settings over available processes and reduce auto-tuning time in the process, and guarantee the same settings across processes.
  2. 2.
    Vast search space. there exist in the order of thousands of valid settings for any given problem size, and searching over all of them is intractable. Our solution so far has been to generate far fewer possible settings, and even reduce the number of settings further by introducing a “thorough mode”, which is disabled by default, but when enabled, will allow users to search over more settings and potentially gain more in speed. This issue is a common problem in modern computational packages, and we hope to alleviate it by common practices such as reducing benchmark time, distributing the process, caching to disk, lazy benchmarking, and approximate performance models.
Appendix B Additional experiments
Herein we present some additional performance metrics from our GEMM-based and fused kernels compared against the baseline.
In Tab. 4, we break down expected performance improvements at the operation level from a single forward and backward pass. Both our GEMM-based and fused kernels provide significant improvement on average over the baseline, while there still exist cases where naive could potentially perform better, especially compared to our GEMM-based kernels. As pointed out in Sec. 3.2, the scatter and gather operation in our GEMM kernels are a significant bottleneck, especially in lower-precision and in NN and IN operations. In lower precision, NN and IN, which account for 75% of the backward pass operations (excluding softmax) will fail to hide their prefetch runtime from global reads, which are not asynchronous, and this will essentially impact the backward pass more than it does the forward pass. This issue, however, is limited to our unfused variant, and our fused kernels maintain their superior performance levels, offering up to an order of magnitude improvement in all variants (1-D, 2-D, and 3-D).
Table 4: Forward + backward pass benchmark breakdown. Improvements over naive, while not as significant as in the forward pass, are still significant. We report benchmark the full forward and backward pass in half precision only, because most training is done in lower precision.
Dim GEMM over naive Fused over naive Fused over GEMM
Average Min Max Average Min Max Average Min Max
1-D $\boldsymbol{\uparrow}$ 502 % $\boldsymbol{\downarrow}$48 % $\boldsymbol{\uparrow}$ 3017 % $\boldsymbol{\uparrow}$ 844 % $\boldsymbol{\downarrow}$20 % $\boldsymbol{\uparrow}$ 7605 % $\boldsymbol{\uparrow}$ 57 % $\boldsymbol{\downarrow}$50 % $\boldsymbol{\uparrow}$ 229 %
[…]